A particle moving along the x-axis in simple harmonic motion starts from the equilibrium position, the origin, at time t=0 and moves to the right. The position of the particle after time, t, is given by the following equation: x=2.0 sin 3πt m



Calculate:



A) the maximum velocity of the motion



B) the acceptation of the object after 2 seconds



C) the distance traveled by the particle from t=0s to t=2s

1
Expert's answer
2021-11-30T17:44:25-0500

a)


"V=2(3\pi)=18.8\frac{m}{s}"

b)


"a=-2(3\pi)^2\sin 3(2)π=0\frac{m}{s^2}"

c)


"X=0\ m"